A 2.2kg mass is whirled in a vertical circle of radius 1.8m at a uniform constant speed. calculate this speed if the minimum tension in the rope is 9.5N.


1
Expert's answer
2021-01-22T17:01:15-0500

"mg+N=m\frac{v^2}{r}\to"


"v=\sqrt{\frac{(N+mg)r}{m}}=\sqrt{\frac{(9.5+2.2\cdot 9.8)\cdot1.8}{2.2}}=5(m/s)" . Answer